What is the electric field at a distance of 10 cm from an isolated point charge of



2 x 10 -9 C?



With solution please?

1
Expert's answer
2022-03-09T12:19:22-0500

The electric field due to the point charge

"E=k\frac{q}{r^2}"

So

"E=9*10^9*\frac{2*10^{-9}}{0.1^2}=1800\:\rm N/C"